# Dynamics of interacting bosons in a two-leg ring ladder with artificial magnetic flux and ac-driven modulations

### v1

@@ Abstract
We investigate the nonequilibrium dynamics of interacting bosons in a two-leg ring ladder pierced by an artificial magnetic flux, where the particles are initially localized in the central sites of both rings, and the ac-driven local energy shifts are applied to the remaining lattice sites. Within the mean-field approximation, we demonstrate the emergence of nonlinear self-trapping for strong interparticle interactions, and characterize the distinct excitation regimes in the absence of the inter-ring tunneling. The artificial magnetic flux typically introduces Peierls phase factors, which induces complex-valued hopping amplitudes and leads to directed net particle currents along the chains. By further incorporating the finite inter-ring coupling and biased intra-ring hopping, we reveal that the tuning of the drive frequency and Peierls phase allows the precise control over both the intensity and direction of particle currents, which facilitates the transition between chiral and antichiral dynamics. These findings offer insights into the coherent manipulation of matter-wave transports in closed-loop lattice configurations and the exploration of nonequilibrium synthetic quantum systems in related fields.
@@ I Introduction
Interacting bosons on a linear ladder in the presence of a uniform magnetic filed has been the topic of previous theoretical works [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], as it is pretty straightforward to study the dynamical responses of particles to magnetic fields, and offer insights into two-dimensional many-body quantum systems. The artificial magnetic flux typically induces phase transitions when the strength of hopping across the ladder and the interactions between the particles are varied, and the specific geometry can be experimentally implemented with ultracold atoms in tunable optical lattices [28, 29, 30, 31, 32, 33, 34, 35], where the chiral dynamics has been observed under precisely adjustable flux and interactions. In addition, it has been demonstrated that the on-site gain and loss in a noninteracting ladder system with the flux-induced asymmetric transport and inter-leg coupling generally results in the antichiral dynamics, where particles in both legs of the ladder propagate unidirectionally in the same direction [36, 37].
@@
There has also been increasing research interest in ultracold quantum gases subjected to the ring-shaped optical lattices [38, 39, 41, 40, 42, 43, 44, 45, 46] and coupled rings [47, 48, 49, 50, 51, 52, 53] due to the promising applications in future atomic circuits, and a finite-length two-leg ladder can readily be mapped to a two-leg ring ladder. Related intriguing physics in this specific configuration has been recently explored [54, 55, 56, 57, 58, 59], while most of the efforts focused on the undriven quantum dynamics or ladders with homogeneous interacting particles on lattice sites. Time-periodic modulation serves as a versatile tool for engineering quantum states and controlling the nonequilibrium dynamics that do not appear in static systems [5], and a combination of the artificial magnetic flux with a joint modulation of both on-site energies and tunneling parameters allows further investigations on persistent currents and quantum simulations in dimensions larger than one. Moreover, the population imbalance between the two rings generates an additional phase, where one may also reorient the direction of the particle currents in the context of an ac-driving and biased intra-ring hopping [7, 45]. In this work, we introduce a two-leg ring ladder pierced by an artificial magnetic flux, where the bosons are initially prepared in the central sites. We apply the ac-driven local energy shifts to the remaining lattice sites, and numerically analyze the excitation regimes for different interactions. We also consider biased intra-ring hopping amplitudes, and modulate the drive frequency and Peierls phase to reveal the chiral and antichiral dynamics.
@@ II Theoretical Model
$+M\left(t\right)\sum_{\ell=1}^{L}\left(\hat{a}_{\ell}^{{\dagger}}\hat{a}_{\ell}+\hat{b}_{\ell}^{{\dagger}}\hat{b}_{\ell}\right)-J_{c}\sum_{\ell=0}^{L}\left(\hat{a}_{\ell}^{{\dagger}}\hat{b}_{\ell}+{\rm H.c.}\right)$
@@
$+M(t)\left[\begin{array}[]{cc}\varphi_{a,\ell}\delta_{\ell,m}\\ \varphi_{b,\ell}\delta_{\ell,m}\end{array}\right],$
@@ III.1 Self-trapping dynamics and excitation regimes
We first analyze the nonlinear interaction effects on the time evolution of particles, as shown in Fig. 2. One can find distinct dynamical evolutions depending on the interaction strength $U$: In the noninteracting case ($U=0$), the particles can readily “escape” from the central site and counterpropagate freely along the chain, resembling a back-and-forth trace at long times with occasional occupation on the lattice sites. At weak interactions (exemplified by $U=2$), the free evolution is slightly weakened, where the particles exhibit enhanced occupation at specific sites over time. As for stronger interactions (e.g., $U=4$ and $U=6$), however, the propagation is significantly suppressed, resulting in the disorder-free localization phenomenon, namely, the self-trapping effect, where the particles remain predominantly trapped in the central site.
@@
The modulation of the local energies with respect to sites $\ell\geq 1$ can be generally regarded as a substrate potential that makes the central site vary from the others, and here we consider the case with the intensity of the shift being constant, i.e., the amplitude of the ac-driven component is $M=0$. As can be plainly seen in Fig. 3, for generic interaction strengths a negative dc amplitude $\mu$ typically strengthens the self-trapping, as it increases the energy offset between the central site and the surrounding sites, thereby the particle number of the central site simply oscillates and the atoms remain bound. Conversely, for moderate positive amplitudes the decay of the particles can be enhanced, with stronger interactions leading to more prominent enhancement, since the effective potential barrier is reduced. In addition, for sufficiently large positive amplitudes $\mu\gg J_{\nu}$ (e.g., $\mu=10$), particle localization can also occur. Note that at $\mu=0$ the system exhibits free evolution, where the decay is the most significant for weak interactions but greatly hindered with rapid oscillation for stronger interactions, which is consistent with Fig. 2. In the following, we primarily take the interaction strength $U=6$ and the dc amplitude $\mu=10$ into account, whence the supercurrents are virtually suppressed before the ac-driven modulation is turned on [45], such that we can explicitly clarify the ac-driven scenario.
Figure 3: Decay of particles in the central site for different dc amplitude $\mu$, when the ac-driving is absent. Here, the interaction strengths are $U=0$, 22, 44 and $6$, respectively, and the Peierls phase is $\theta=0$. The inter-ring coupling strength is $J_{c}=0$, and the intra-ring hopping amplitude is $J_{\nu}=1$.
With the application of the ac-driving, the local energies of sites $\ell\geq 1$ are shifted periodically in time. One can thus define the difference of the particle number
@@
Here we are mainly interested in the short-time behaviors before the outgoing particles return to the central site, so we perform the simulation over $0<t_{s}<50$, and calculate the maximum difference $|\Delta_{\nu,0}|_{\rm{max}}=\underset{t\in t_{s}}{\rm{max}}|\Delta_{\nu,0}\left(t\right)|$. Figure 4 illustrates how the maximum difference depends on the drive strength $M$ and the drive frequency $\omega$. For generic drive frequencies and weak ac-driven modulation amplitude, the system can be quite stable and only a tiny portion of particles are excited. We see that for roughly $2J_{\nu}<\omega-\Delta\epsilon<2J_{\nu}$, with $\Delta\epsilon$ denoting the energy gap between the central site and the nearest-neighboring sites, even a small $M$ leads to particle excitation. This is related to the energy conservation: In the noninteracting limit $U=0$, one can directly perform the Fourier transform on Hamiltonian (1) to obtain the Bloch Hamiltonian $H\left(k\right)$ in the momentum space, and figure out the dispersion relation $\epsilon_{k}=-2J_{\nu}\cos\left(k\right)$, where in our case the gap is explicitly the dc amplitude $\mu$, requiring the drive frequency $2J_{\nu}+\mu<\omega<2J_{\nu}+\mu$. At weak interactions, in combination with the Bogoliubov approximation [61] gives the gap
@@
where finite interaction strengths $U$ specifically renormalize the boundaries of the bands. There are further bands of instability with $2J_{\nu}+\Delta\epsilon<n\omega<2J_{\nu}+\Delta\epsilon$ for integers $n$, as the drive also provides energy in multiples, and the $n$th order bands overlap with each other. Note that for the drive frequency $\omega\approx 0$, the ac-driven component $M\cos\left(\omega t\right)$ can be readily absorbed into the dc amplitude $\mu$ as an even larger positive amplitude, and hence few particles escape from the central site.
Figure 4: The maximum particle number difference $|\Delta_{\nu,0}|_{\rm{max}}$ vs drive strength $M$ and drive frequency $\omega$ under different interaction strength $U=0$, 22, 44, and $6$, respectively. Here, we have taken the dc amplitude $\mu=10$, the intra-ring hopping amplitude $J_{\nu}=1$, the inter-ring coupling strength $J_{c}=0$, and the Peierls phase $\theta=0$. The simulation time is $t_{s}=50$, and the dashed lines delineate the regimes $\omega_{\pm}=\left(\pm 2J_{\nu}+\Delta\epsilon\right)/n$ corresponding to the $n$th order excitations.
The Peierls phase typically breaks the time-invariance symmetry, where the particles hop on the lattice with complex tunneling matrix elements. Basically, the particle current flowing from site $\ell$ to $\ell+1$ in ring $\nu$ is written as $I_{\nu,\ell}\left(t\right)=\langle-\partial\hat{H}/\partial\theta\rangle$, which yields
@@
As shown in Fig. 5, when the ac-driving is applied, the net particle currents for the phases $\theta=\pi/8$ and $\theta=\pi/4$ exhibit rapid oscillations around typical “negative” values (clockwise flows based on the sketch of Fig. 1), with their amplitudes slightly damping with time. In sharp contrast, for $\theta=5\pi/8$ and $\theta=3\pi/4$ both currents reverse directions, oscillating around “positive” values while maintaining the same magnitudes as their counterparts ($\theta=\pi/8$ and $\theta=\pi/4$, respectively). This correspondence arises from the equivalencies of the exponential, where $e^{i\left(\alpha+\pi/2\right)}=ie^{i\alpha}$ ($\alpha$ is real) and the currents obey $I_{\nu}(t)_{\theta=\alpha}=-I_{\nu}(t)_{\theta=\alpha+\pi/2}$, and at the specific phase $\theta=\pi/2$ the current appears to oscillate around zero. As a consequence, the direction of $I_{\nu}$ can be simply varied by either inverting the Peierls phase from $\theta$ to $\theta$, or making use of the phase-dependent symmetry.
@@
There are also two exemplified time scales $\tau_{-}$ and $\tau_{+}$ illustrated in Fig. 5. Since the ac-driving follows a cosinusoidal oscillation, the total amplitude stays even larger than the fixed dc amplitude during the typical half cycle ($t\in\left[\tau_{-},\tau_{+}\right]$), which is followed by a reciprocal tendency, i.e., its intensity first increases and then decreases, leading to the “revival” of the particle current. In the other half of the driving cycle, however, the total amplitude keeps smaller than the dc amplitude, which stimulates the excitations with a relatively smoother oscillation, and results in a “comb-like” pattern overall. Notably, at sufficiently long times the particle occupation on the lattice sites eventually saturates, and the system reaches a dynamic equilibrium, where the currents along the chain remain small but nearly persistent.
Figure 6: The maximum particle imbalance $|\Delta N|_{\rm{max}}$ as a function of the dc amplitude $\mu$ for different hopping amplitude $J_{b}$. Here, the inter-ring coupling strength is $J_{c}=1$, and the interaction strength is $U=6$. The ac-driving is absent, and the calculation is performed with $0<t_{s}<50$.
@@ III.2 Chiral and antichiral dynamics with biased intra-ring hopping
We now turn to study the dynamics when the inter-ring coupling strength $J_{c}$ is nonzero, where the corresponding lattices are coupled and the system changes into a “two-leg ring ladder”. For convenience, we keep $J_{c}=1$, and still lay emphasis on the interaction strength $U=6$ in order to maintain consistency with the former case. The particle imbalance between the rings can be defined as
@@
The influence of the drive frequency $\omega$ on $\overline{\mathcal{M}}_{\nu}$ can also be examined, as illustrated in Fig. 7(b). For a driving interval of $\tau=50$, one can find $\overline{\mathcal{M}}_{a}\approx\overline{\mathcal{M}}_{b}\approx 0$ at $\omega=0$, since most of the particles are still trapped in the central sites of both rings under such conditions. At around $\omega=1$, $\overline{\mathcal{M}}_{\nu}$ peaks with $\overline{\mathcal{M}}_{a}>\overline{\mathcal{M}}_{b}$, and beyond this point they decrease while oscillating alternately with the increase of the drive frequency $\omega$, corresponding to the particle exchange between the rings. For $\omega>10$, they stabilize with $\overline{\mathcal{M}}_{a}\gtrsim\overline{\mathcal{M}}_{b}$ even when the drive frequency is further increased, which is consistent with the excitation regimes of the system. In the other case of $\tau=100$, the behaviors of $\overline{\mathcal{M}}_{a}$ and $\overline{\mathcal{M}}_{b}$ roughly mirror their counterparts at $\tau=50$, while maintaining smaller magnitudes overall. Note that in the ring-lattice geometry, particles undergo back-and-forth motions (c.f. Fig. 2), and for sufficiently long driving intervals the occupation on specific sites may gradually saturate, which differs from a two-leg ladder that the population stacks at one end manifesting the skin dynamics [24].
@@
We further take the moderate bias $\delta J=0.6$ as a typical example, and illustrate the dependence of the mean particle currents on the drive frequency $\omega$ and the Peierls phase $\theta$. The asymptotic mean particle current along ring $\nu$ can be calculated by $\bar{I}_{\nu}=\langle I_{\nu}\left(t\right)\rangle_{\tau}$, where $\langle\ldots\rangle_{\tau}$ represents the time average over a driving interval $\tau$. As can be seen in Fig. 8, at low frequency $\omega=1$ currents $\bar{I}_{a}$ and $\bar{I}_{b}$ maintain identical directions with different intensities, while they both reverse when the drive frequency is tuned to $\omega=3$, with one of the intensities being increased and the other being decreased. For the case of $\omega=4$, the two currents flow in opposite directions, and at $\omega=7$ this pattern continues, which roughly mirrors the behavior of $\omega=4$, with the intensity of $\bar{I}_{a}$ being almost unchanged and that of $\bar{I}_{b}$ being slightly reduced. The currents realign at higher frequencies $\omega=9$ and $\omega=11$, where they again flow in the same direction.
@@
Figure 9: Regimes of the chiral and antichiral dynamics for interaction strength $U=0$ and $U=6$, respectively. (a) and (c): $\bar{I}_{a}\bar{I}_{b}$ vs the drive frequency $\omega$ and the Peierls phase $\theta$. (b) and (d): The phase diagram in the $\theta$-$\omega$ plane. The blue shading denotes the region for the chiral dynamics, while the orange shading corresponds to that of the antichiral dynamics. Here, we have taken $J_{c}=1$, $\delta J=0.6$, $\mu=10$, and $M=8$, and the driving interval is kept as $\tau=10$.
@@
To explore more quantitatively the chiral and antichiral dynamics, we identify the directions of the particle currents along the two rings under different interactions, where they have opposite directions with $\bar{I}_{a}\bar{I}_{b}<0$ and the same direction with $\bar{I}_{a}\bar{I}_{b}>0$, respectively, which can be visible after short-time evolutions. One can see from Figs. 9(a) and 9(c) that the introduction of the particle interaction and the ac-driven modulation notably shifts the bands with increased overlap regions, and also varies the relative directions of particle currents between the rings. At the interaction strength $U=6$, there are substantial regions corresponding to $\bar{I}_{a}\bar{I}_{b}<0$, where the chiral dynamics is exhibited, while at $U=0$ the value of $\bar{I}_{a}\bar{I}_{b}$ keeps predominantly positive, showcasing explicitly the antichiral dynamics. The phase $\theta=\pi/2$ serves as the cental line for the excitations according to the exponential equivalencies, with the features repeating on both sides. We characterize the chiral (blue shading) and antichiral (orange shading) dynamics in the $\theta$-$\omega$ plane by analyzing the sign of $\bar{I}_{a}\bar{I}_{b}$ in Figs. 9(b) and 9(d), where $\bar{I}_{a}\bar{I}_{b}\approx 0$ marks the chiral-antichiral transition boundaries. In the noninteracting limit ($U=0$), only tiny discrete regions exhibit chiral dynamics, while at $U=6$ the transition points shift significantly, and the chiral regions expand while the antichiral regions contract. This trend suggests that with further increases in the interaction strength, the antichiral dynamics would be suppressed, and the transition disappears ultimately due to the self-trapping effect. Accordingly, the intensity and direction of particle currents can be systematically controlled by manipulating the chiral-anticiral transition through the tunings of the drive frequency and Peierls phase.

### v2

@@ Abstract
We investigate the nonequilibrium dynamics of interacting bosons in a two-leg ring ladder pierced by an artificial magnetic flux, where the particles are initially localized in the central sites of both rings, and the ac-driven local energy shifts are applied to the remaining lattice sites. Within the mean-field approximation, we demonstrate the emergence of nonlinear self-trapping for strong interparticle interactions, and characterize the distinct excitation regimes in the absence of the inter-ring tunneling. The artificial magnetic flux typically introduces the Peierls phase factors, which induces complex-valued hopping amplitudes and leads to directed net particle currents along the chains. By further incorporating the finite inter-ring coupling and biased intra-ring hopping, we reveal that the tuning of the drive frequency and Peierls phase allows the precise control over both the intensity and direction of particle currents, which facilitates the transition between chiral and antichiral dynamics. These findings offer insights into the coherent manipulation of matter-wave transports in closed-loop lattice configurations and the exploration of nonequilibrium synthetic quantum systems in related fields.
@@ I Introduction
Interacting bosons on a linear ladder in the presence of a uniform magnetic filed has been the topic of previous theoretical works [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], as it is pretty straightforward to study the dynamical responses of particles to magnetic fields, and offer insights into two-dimensional many-body quantum systems. The artificial magnetic flux typically induces phase transitions when the strength of hopping across the ladder and the interactions between the particles are varied, and the specific geometry can be experimentally implemented with ultracold atoms in tunable optical lattices [30, 31, 32, 33, 34, 35, 36, 37], where the chiral dynamics has been observed under precisely adjustable flux and interactions. In addition, it has been demonstrated that the on-site gain and loss in a noninteracting ladder system with the flux-induced asymmetric transport and inter-leg coupling generally results in the antichiral dynamics, where particles in both legs of the ladder propagate unidirectionally in the same direction [38, 39].
@@
There has also been increasing research interest in ultracold quantum gases subjected to the ring-shaped optical lattices [40, 41, 42, 43, 44, 45, 46, 47, 48] and coupled rings [49, 50, 51, 52, 53, 54, 55] due to the promising applications in future atomic circuits, and a finite-length two-leg ladder can readily be mapped to a two-leg ring ladder. Related intriguing physics in this specific configuration has been recently explored [56, 57, 58, 59, 60, 61], while most of the efforts focused on the undriven quantum dynamics or ladders with homogeneous interacting particles on lattice sites. Time-periodic modulation serves as a versatile tool for engineering quantum states and controlling the nonequilibrium dynamics that do not appear in static systems [5], and a combination of the artificial magnetic flux with a joint modulation of both on-site energies and tunneling parameters allows further investigations on persistent currents and quantum simulations in dimensions larger than one. Moreover, the population imbalance between the two rings generates an additional phase, where one may also reorient the direction of the particle currents in the context of an ac driving and biased intra-ring hopping [7, 47]. In this work, we introduce a two-leg ring ladder pierced by an artificial magnetic flux, where the bosons are initially prepared in the central sites. We apply the ac-driven local energy shifts to the remaining lattice sites, and numerically analyze the excitation regimes for different interactions. We also consider biased intra-ring hopping amplitudes, and modulate the drive frequency and Peierls phase to reveal the chiral and antichiral dynamics.
@@ II Theoretical Model
$+F\left(t\right)\sum_{\ell=1}^{L}\left(\hat{a}_{\ell}^{{\dagger}}\hat{a}_{\ell}+\hat{b}_{\ell}^{{\dagger}}\hat{b}_{\ell}\right)-J_{c}\sum_{\ell=0}^{L}\left(\hat{a}_{\ell}^{{\dagger}}\hat{b}_{\ell}+{\rm H.c.}\right)$
@@
$+F(t)\left[\begin{array}[]{cc}\varphi_{a,\ell}\delta_{\ell,m}\\ \varphi_{b,\ell}\delta_{\ell,m}\end{array}\right],$
@@ III.1 Self-trapping dynamics and excitation regimes
We first analyze the nonlinear interaction effects on the time evolution of particles, as shown in Fig. 2. One can find distinct dynamical evolutions depending on the interaction strength $U$: in the noninteracting case ($U=0$), the particles can readily escape from the central site and counterpropagate freely along the chain, resembling a back-and-forth trace at long times with occasional occupation on the lattice sites. At weak interactions (exemplified by $U=2$), the free evolution is slightly weakened, where the particles exhibit enhanced occupation at specific sites over time. As for stronger interactions (e.g., $U=4$ and $U=6$), however, the propagation is significantly suppressed, resulting in the disorder-free localization phenomenon, namely, the self-trapping effect, where the particles remain predominantly confined in the central site.
@@
The modulation of the local energies with respect to sites $\ell\geq 1$ can be generally regarded as a substrate potential that makes the central site vary from the others, and here we consider the case with the intensity of the shift being constant, i.e., the amplitude of the ac-driven component is $M=0$. As can be plainly seen in Fig. 3, for generic interaction strengths a negative dc amplitude $\mu$ typically strengthens the self-trapping, as it increases the energy offset between the central site and the surrounding sites, thereby the particle number of the central site simply oscillates and the atoms remain bound. Conversely, for moderate positive amplitudes the decay of the particles can be enhanced, with stronger interactions leading to more prominent enhancement and oscillations, since in this case the central site may be intuitively considered as a potential well, and the effective potential difference holds modest. In addition, for sufficiently large positive amplitudes $\mu\gg J_{\nu}$ (e.g., $\mu=10$), particle localization can also occur. Note that at $\mu=0$ the system exhibits free evolution, where the decay is the most significant for weak interactions but greatly hindered with rapid oscillation for stronger interactions, which is consistent with Fig. 2. In what follows, we primarily take the dc amplitude $\mu=10$ into account, whence the supercurrents are virtually suppressed before the ac-driven modulation is turned on [47], such that we can explicitly clarify the ac-driven scenario.
Figure 3: Decay of particles in the central site for different dc amplitude $\mu$, when the ac driving is absent. Here, the interaction strengths are $U=0$, 22, 44 and $6$, respectively, and the Peierls phase is $\theta=0$. The inter-ring coupling strength is $J_{c}=0$, and the intra-ring hopping amplitude is $J_{\nu}=1$.
With the application of the ac driving, the local energies of sites $\ell\geq 1$ are shifted periodically in time. One can thus define the difference of the particle number
@@
Here we are mainly interested in the short-time behaviors before the outgoing particles return to the central site, so we perform the simulation over $0<t_{s}<50$, and calculate the maximum difference $|\Delta_{\nu,0}|_{\rm{max}}=\underset{t\in t_{s}}{\rm{max}}|\Delta_{\nu,0}\left(t\right)|$. Figure 4 illustrates how the maximum difference depends on the drive strength $M$ and the drive frequency $\omega$. For generic drive frequencies and weak ac-driven modulation amplitude, the system can be quite stable and only a tiny portion of particles are excited. We see that for roughly $2J_{\nu}<\omega-\Delta\epsilon<2J_{\nu}$, with $\Delta\epsilon$ denoting the energy gap between the central site and the nearest-neighboring sites, even a small $M$ leads to particle excitation. This is related to the energy conservation: In the noninteracting limit $U=0$ and in the absence of the Peierls phase and ac modulation, one can directly perform the Fourier transform on Hamiltonian (1) to obtain the Bloch Hamiltonian $H\left(k\right)$ in the momentum space and figure out the dispersion relation $\epsilon_{k}=-2J_{\nu}\cos\left(k\right)$, where in our case the gap is explicitly the dc amplitude $\mu$, requiring the drive frequency $2J_{\nu}+\mu<\omega<2J_{\nu}+\mu$ for exciting the particles. At weak interactions, in combination with the Bogoliubov approximation [63] gives the gap
@@
where finite interaction strengths $U$ specifically renormalize the boundaries of the bands. There are further bands of instability with $2J_{\nu}+\Delta\epsilon<n\omega<2J_{\nu}+\Delta\epsilon$ for integers $n$, as the drive also provides energy in multiples, and the $n$th-order bands overlap with each other. Within the allowed regions, the ac amplitude mainly affects the intensity of the tunneling, where a larger $M$ typically results in more excited particles with a larger $|\Delta_{\nu,0}|_{\rm{max}}$. Note that for the drive frequency $\omega\approx 0$, the ac-driven component $M\cos\left(\omega t\right)$ can be readily absorbed into the dc amplitude $\mu$ as an even larger positive amplitude, and hence few particles escape from the central site.
Figure 4: The maximum particle number difference $|\Delta_{\nu,0}|_{\rm{max}}$ vs drive strength $M$ and drive frequency $\omega$ under different interaction strength $U=0$, 22, 44, and $6$, respectively. Here, we have taken the dc amplitude $\mu=10$, the intra-ring hopping amplitude $J_{\nu}=1$, the inter-ring coupling strength $J_{c}=0$, and the Peierls phase $\theta=0$. The simulation time is $t_{s}=50$, and the dashed lines delineate the regimes $\omega_{\pm}=\left(\pm 2J_{\nu}+\Delta\epsilon\right)/n$ corresponding to the $n$th-order excitations.
The Peierls phase typically breaks the time-reversal symmetry, where the particles hop on the lattice with complex tunneling matrix elements. Basically, the particle current flowing from site $\ell$ to $\ell+1$ in ring $\nu$ yields
@@
As shown in Fig. 5, when the ac driving is applied, the net particle currents for the phases $\theta=\pi/8$ and $\theta=\pi/4$ exhibit rapid oscillations around typical negative values (clockwise flows based on the sketch of Fig. 1), with their amplitudes slightly damping with time. In sharp contrast, for $\theta=5\pi/8$ and $\theta=3\pi/4$ both currents reverse directions, oscillating around positive values while maintaining the same magnitudes as their counterparts ($\theta=\pi/8$ and $\theta=\pi/4$, respectively). This correspondence arises from the equivalencies of the exponential, where $e^{i\left(\alpha+\pi/2\right)}=ie^{i\alpha}$ ($\alpha$ is real) and the currents obey $I_{\nu}(t)_{\theta=\alpha}=-I_{\nu}(t)_{\theta=\alpha+\pi/2}$, and at the specific phase $\theta=\pi/2$ the current appears to oscillate around zero. As a consequence, the direction of $I_{\nu}$ can be simply varied by either inverting the Peierls phase from $\theta$ to $\theta$, or making use of the phase-dependent symmetry.
@@
There are also two exemplified time scales $\tau_{-}$ and $\tau_{+}$ illustrated in Fig. 5. Since the ac driving follows a cosinusoidal oscillation, the total amplitude stays even larger than the fixed dc amplitude during the typical half cycle ($t\in\left[\tau_{-},\tau_{+}\right]$), which is followed by a reciprocal tendency, i.e., its intensity first increases and then decreases, leading to the revival of the particle current. In the other half of the driving cycle, however, the total amplitude keeps smaller than the dc amplitude, which stimulates the excitations with a relatively smoother oscillation, and results in a comblike pattern overall. Notably, at sufficiently long times the particle occupation on the lattice sites eventually saturates, and the system reaches a dynamic equilibrium, where the currents along the chain remain small but nearly persistent.
Figure 6: The maximum particle imbalance $|\Delta N|_{\rm{max}}$ as a function of the dc amplitude $\mu$ for different hopping amplitude $J_{b}$. Here, the inter-ring coupling strength is $J_{c}=1$, and the interaction strength is $U=6$. The ac driving is absent, and the calculation is performed with $0<t_{s}<50$.
@@ III.2 Chiral and antichiral dynamics with biased intra-ring hopping
We now turn to study the dynamics when the inter-ring coupling strength $J_{c}$ is nonzero, where the corresponding lattices are coupled and the system changes into a two-leg ring ladder. For convenience, we keep $J_{c}=1$, and still lay emphasis on the interaction strength $U=6$ in order to maintain consistency with the former case. The particle imbalance between the rings can be defined as
@@
The influence of the drive frequency $\omega$ on $\overline{\mathcal{M}}_{\nu}$ can also be examined, as illustrated in Fig. 7(b). For a driving interval of $\tau=50$, one can find $\overline{\mathcal{M}}_{a}\approx\overline{\mathcal{M}}_{b}\approx 0$ at $\omega=0$, since most of the particles are still confined in the central sites of both rings under such conditions. At around $\omega=1$, $\overline{\mathcal{M}}_{\nu}$ peaks with $\overline{\mathcal{M}}_{a}>\overline{\mathcal{M}}_{b}$, and beyond this point they decrease while oscillating alternately with the increase of the drive frequency $\omega$, corresponding to the particle exchange between the rings. For $\omega>10$, they stabilize with $\overline{\mathcal{M}}_{a}\gtrsim\overline{\mathcal{M}}_{b}$ even when the drive frequency is further increased, which is consistent with the excitation regimes of the system. In the other case of $\tau=100$, the behaviors of $\overline{\mathcal{M}}_{a}$ and $\overline{\mathcal{M}}_{b}$ roughly mirror their counterparts at $\tau=50$, while maintaining smaller magnitudes overall. Note that in the ring-lattice geometry, particles undergo back-and-forth motions (c.f. Fig. 2), and for sufficiently long driving intervals the occupation on specific sites may gradually saturate, which differs from a two-leg ladder that the population stacks at one end manifesting the skin dynamics [24].
@@
We further take the moderate bias $\delta J=0.6$ as a typical example, and illustrate the dependence of the mean particle currents on the drive frequency $\omega$ and the Peierls phase $\theta$. The asymptotic mean particle current along ring $\nu$ can be calculated by $\bar{I}_{\nu}=\langle I_{\nu}\left(t\right)\rangle_{\tau}$, where $\langle\ldots\rangle_{\tau}$ represents the time average over a driving interval $\tau$. As can be seen in Fig. 8, at low frequency $\omega=1$ currents $\bar{I}_{a}$ and $\bar{I}_{b}$ maintain identical directions with different intensities, while they both reverse when the drive frequency is tuned to $\omega=3$, with one of the intensities being increased and the other being decreased. For the case of $\omega=4$, the two currents flow in opposite directions, and at $\omega=7$ this pattern continues, which roughly mirrors the behavior of $\omega=4$, and the intensity of $\bar{I}_{a}$ keeps almost unchanged and that of $\bar{I}_{b}$ is slightly reduced. The currents realign at higher frequencies $\omega=9$ and $\omega=11$, where they again flow in the same direction.
@@
Figure 9: Regimes of the chiral and antichiral dynamics for interaction strength $U=0$ and $U=6$, respectively. (a) and (c) $\bar{I}_{a}\bar{I}_{b}$ vs the drive frequency $\omega$ and the Peierls phase $\theta$. (b) and (d) The phase diagram in the $\theta$-$\omega$ plane. The blue shading denotes the region for the chiral dynamics, while the orange shading corresponds to that of the antichiral dynamics. Here, we have taken $J_{c}=1$, $\delta J=0.6$, $\mu=10$, and $M=8$, and the driving interval is kept as $\tau=10$.
@@
To explore more quantitatively the chiral and antichiral dynamics, we identify the directions of the particle currents along the two rings under different interactions, where they have opposite directions with $\bar{I}_{a}\bar{I}_{b}<0$ and the same direction with $\bar{I}_{a}\bar{I}_{b}>0$, respectively, which can be visible after short-time evolutions. One can see from Figs. 9(a) and 9(c) that the introduction of the particle interaction and the ac-driven modulation notably shifts the bands with increased overlap regions, and also varies the relative directions of particle currents between the rings. For the interaction strength $U=6$, there are substantial regions corresponding to $\bar{I}_{a}\bar{I}_{b}<0$, where the chiral dynamics is exhibited, while at $U=0$ the value of $\bar{I}_{a}\bar{I}_{b}$ keeps predominantly positive, showcasing explicitly the antichiral dynamics. The phase $\theta=\pi/2$ serves as the cental line for the excitations according to the exponential equivalencies, with the features repeating on both sides. We characterize the chiral (blue shading) and antichiral (orange shading) dynamics in the $\theta$-$\omega$ plane by analyzing the sign of $\bar{I}_{a}\bar{I}_{b}$ in Figs. 9(b) and 9(d), where $\bar{I}_{a}\bar{I}_{b}\approx 0$ marks the chiral-antichiral transition boundaries. In the noninteracting case, the chirality appears at well-defined, discrete narrow bands in frequency $\omega$, and the spectrum is antisymmetric around $\theta=\pi/2$, while at $U=6$ the bands are significantly enlarged and fragmented for broader ranges of frequency, and thus the antichiral regions relatively contract. This suggests that chirality survives strong interactions, and their deformation signals a transition between the chiral and antichiral dynamics, which disappears ultimately due to the self-trapping effect. Accordingly, the intensity and direction of particle currents can be systematically controlled by manipulating the chiral-anticiral transition through joint tunings of the drive frequency and the Peierls phase.